# One Pixel is All I Need

### v1

@@ Abstract
Since their introduction, Vision Transformers (ViTs) have demonstrated record-breaking performance in various visual tasks. However, their robustness against potential malicious attacks has also attracted increasing attention from researchers. Backdoor attacks involve forming a strong association between a specific trigger and a target label. The backdoored model correctly identifies clean images but predicts the attacker-specified label when the trigger is present in the input. Previous work has attempted to use triggers that are different but similar to those used during training. We refer to such patterns that can still trigger the backdoor despite being different from the original trigger as quasi-triggers. We found that, compared to CNNs, ViTs exhibit higher attack success rates for quasi-triggers. Additionally, when some backdoor features are present in clean samples, the backdoor may be suppressed in some inputs, causing the original trigger to fail, while quasi-triggers can successfully trigger the backdoor and complete the attack.
The perturbation sensitivity distribution map (PSDM) of the model is generated by computing and summing gradients over a large number of inputs. This map shows the model’s sensitivity to content-agnostic perturbations in the input. Utilizing the PSDM can help us better alter the model’s decisions, an angle that past malicious attacks have not considered. Notably, the PSDM of ViTs exhibits a clear patch-like pattern, where the central pixels of each patch are more sensitive to perturbations than the edges. The PSDM is specifically used to guide the use of quasi-triggers.
Based on these findings, we designed a simple yet effective data poisoning backdoor and its variants for ViT models. We only need an extremely low poisoning rate, train for one epoch, and modify a single pixel to successfully attack all validation images. We name this attack ”WorstVIT.”
@@ 4.2 Probing Attacks
Coordinates with High PSDM Values: Quasi-triggers are sequentially placed at coordinates in the Perturbation Sensitivity Distribution Map (PSDM) where the values are high. Pixels with High Gradient Values: Quasi-triggers are placed at pixel coordinates of the current input image where the gradient values are high. Channels with High Gradient Values: Quasi-triggers are placed at coordinates corresponding to channels with high gradient values. After placing the quasi-triggers, their pixel values are adjusted to optimize attack effectiveness. Once a successful attack is achieved, further fine-tuning of the pixel values continues, selecting quasi-triggers that not only successfully activate the backdoor but also remain visually inconspicuous for each specific sample.
@@ 6.5 WorstVIT and Its Variants
Probing Attack: In the basic assumption of backdoor attacks, the attacker often has full knowledge of the model’s architecture and parameters. This means that before launching a formal attack, we test the feasibility of the attack locally. We select the coordinates of the top 1000 values in the model’s Perturbation Sensitivity Distribution Map (PSDM) and the coordinates of the top 1000 pixels or channels with the highest gradients in the current input as candidate coordinates. We iterate through these coordinates to select those that can successfully attack and remain sufficiently hidden. In a full-to-full attack, we can modify multiple pixels within the same sample until the attack succeeds, which typically requires fewer than 1000 attempts.
Single-Channel Attack: By changing only one channel value, we achieved an attack success rate of 61.21%. Due to the high computational cost, we did not conduct single-channel probing attacks.
all-to-all Attack: As shown in Figure 9, we found that the samples successfully attacked highly overlap with those predicted correctly. This suggests that the model has established a mapping relationship between the true labels and the backdoor labels. Therefore, it is necessary to first predict the correct clean sample before modifying the label to the correct backdoor label. However, we can still modify the labels to our desired results using quasi-triggers and probing attacks.
@@ 6.7 Backdoor Defense
Fine-Pruning: We applied Fine-Pruning to the model as described in the paper, with pruning rates ranging from 0.1 to 0.9. We observed the clean performance and attack success rate before and after pruning. Both metrics decreased almost synchronously, indicating that Fine-Pruning is ineffective against our attack.
@@
SCALE UP: We used SCALE UP with thresholds ranging from 0.1 to 0.9, increasing by 0.2 each time. SCALE UP showed minimal difference in distinguishing between clean and poisoned inputs, making it unable to identify backdoor samples.
@@
Neural Cleanse: Given the large number of classes in ImageNet, we applied Neural Cleanse to downstream tasks for simplicity. We detected the Mean Absolute Deviation (MAD) for labels 0-9 in each downstream task, with the backdoor label being 5. We attempted to reverse-engineer the trigger and compute the MAD with and without regularization of the mask matrix, but both attempts failed. Neural Cleanse was unable to reverse-engineer the trigger, detect the presence of a backdoor, or correctly label the samples if a backdoor was detected.

### v2

@@ Abstract
Since their introduction, Vision Transformers (ViTs) have demonstrated record-breaking performance in various visual tasks. However, their robustness against potential malicious attacks has also attracted increasing attention from researchers. Backdoor attacks involve forming a strong association between a specific trigger and a target label. The backdoored model correctly identifies clean images but predicts the attacker-specified label when the trigger is present in the input. Previous work has attempted to use triggers that are different but similar to those used during training. We refer to such patterns that can still trigger the backdoor despite being different from the original trigger as quasi-triggers. We found that, compared to CNNs, ViTs exhibit higher attack success rates for quasi-triggers. Additionally, when some backdoor features are present in clean samples, the backdoor may be suppressed in some inputs, causing the original trigger to fail, while quasi-triggers can successfully trigger the backdoor and complete the attack. The perturbation sensitivity distribution map (PSDM) of the model is generated by computing and summing gradients over a large number of inputs. This map shows the model’s sensitivity to content-agnostic perturbations in the input. Utilizing the PSDM can help us better alter the model’s decisions, an angle that past malicious attacks have not considered. Notably, the PSDM of ViTs exhibits a clear patch-like pattern, where the central pixels of each patch are more sensitive to perturbations than the edges. The PSDM is specifically used to guide the use of quasi-triggers. Based on these findings, we designed a simple yet effective data poisoning backdoor and its variants for ViT models. We only need an extremely low poisoning rate, train for one epoch, and modify a single pixel to successfully attack all validation images. We name this attack ”WorstVIT.”
@@ 4.2 Probing Attacks
Coordinates with High PSDM Values: Quasi-triggers are sequentially placed at coordinates in the Perturbation Sensitivity Distribution Map (PSDM) where the values are high. Pixels with High Gradient Values: Quasi-triggers are placed at pixel coordinates of the current input image where the gradient values are high. Channels with High Gradient Values: Quasi-triggers are placed at coordinates corresponding to channels with high gradient values. In a full-to-full attack, we can modify multiple pixels within the same sample until the attack succeeds. After placing the quasi-triggers, their pixel values are adjusted to optimize attack effectiveness. Once a successful attack is achieved, further fine-tuning of the pixel values continues, selecting quasi-triggers that not only successfully activate the backdoor but also remain visually inconspicuous for each specific sample.
@@ 6.5 WorstVIT and Its Variants
Hidden-WorstVIT. In the Hidden-WorstVIT variant, during the training phase, we restrict the trigger’s position to one of the four corners of each patch. In the attack phase, this approach remains consistent with WorstVIT. After training is complete, due to the lower sensitivity of the Vision Transformer (ViT) model at patch edges, the poisoned samples in the training set do not exhibit excessive toxicity. This helps evade various detection mechanisms. Moreover, under this poisoning strategy, the attack success rate in the central regions of the patches remains significantly higher compared to the edge regions. This characteristic ensures that the effectiveness of the backdoor attack is maintained while minimizing the risk of detection.
WorstSwin-VIT. In the WorstSwin-VIT variant, as illustrated in Figure 9, we randomly select 1,000 images from the CIFAR-10 dataset and compute the PSDM for the Swin Transformer. The PSDM is divided into multiple $4\times 4$ patch regions. For each patch, the central region exhibits higher PSDM values compared to the peripheral areas, with the highest value observed at position $(108,209)$. Therefore, during the training process of WorstSwin-VIT, instead of dividing images into $16\times 16$ patches, we randomly select a point in the image and modify the pixel values within its surrounding $3\times 3$ neighborhood (i.e., the selected point plus its eight neighboring pixels). During the testing phase, the trigger is consistently placed at the coordinate $(108,209)$.
Figure 9: PSDM of Swin Base Patch4 Window7 224.
Single-Channel Attack. By changing only one channel value, we achieved an attack success rate of 61.21%. Due to the high computational cost, we did not conduct single-channel probing attacks.
all-to-all Attack. As shown in Figure 10, we found that the samples successfully attacked highly overlap with those predicted correctly. This suggests that the model has established a mapping relationship between the true labels and the backdoor labels. Therefore, it is necessary to first predict the correct clean sample before modifying the label to the correct backdoor label. However, we can still modify the labels to our desired results using quasi-triggers and probing attacks.
@@ 6.7 Backdoor Defense
Fine-Pruning. We applied Fine-Pruning to the model as described in the paper, with pruning rates ranging from 0.1 to 0.9. We observed the clean performance and attack success rate before and after pruning. Both metrics decreased almost synchronously, indicating that Fine-Pruning is ineffective against our attack.
@@
SCALE UP. We used SCALE UP with thresholds ranging from 0.1 to 0.9, increasing by 0.2 each time. SCALE UP showed minimal difference in distinguishing between clean and poisoned inputs, making it unable to identify backdoor samples.
@@
Neural Cleanse. Given the large number of classes in ImageNet, we applied Neural Cleanse to downstream tasks for simplicity. We detected the Mean Absolute Deviation (MAD) for labels 0-9 in each downstream task, with the backdoor label being 5. We attempted to reverse-engineer the trigger and compute the MAD with and without regularization of the mask matrix, but both attempts failed. Neural Cleanse was unable to reverse-engineer the trigger, detect the presence of a backdoor, or correctly label the samples if a backdoor was detected.